OpenStudy (anonymous):

12/18 = what over 27

OpenStudy (anonymous):

@calculusxy

OpenStudy (anonymous):

do you not know how to use a calculator?

OpenStudy (anonymous):

I guess I'm dumb

OpenStudy (anonymous):

12/18 = x/27 (cross multiply) 324 = 18x x=18

OpenStudy (anonymous):

hold on don't tell me

OpenStudy (anonymous):

well were all dumb when it comes down to something

OpenStudy (anonymous):

ok, I still don't get it

OpenStudy (anonymous):

I tried to reduce the 1st fraction but I don't know, I just can't think of what it could be.

OpenStudy (anonymous):

oh I think it's 18

OpenStudy (anonymous):

12/18 = x/27 so we are finding x 12/18 x 27 = x 324/18 = x 18=x

OpenStudy (anonymous):

I took 12/18 divided them by two on both sides got 6/9 times 3 on both sides and got 18/27, so I believe 12/18 equals 18/27 right?

OpenStudy (anonymous):

let x is over 27 hence \[\huge \frac{12}{18} = \frac{x}{27} \] let use cross multiplication to solve it \[12 \times 27 = 18 \times x \rightarrow \huge x = \frac{12 \times 27}{18} \] \[\huge x = \frac{2 \times 27}{3} = \frac{2 \times 9}{1} = 18\] so 18 is over 27

OpenStudy (anonymous):

yay! I knew it, so I'm not so dumb after all.. lol